# Co-evolution of Nuclear Star Clusters and Massive Black Holes: Extreme Mass-Ratio Inspirals

### v1

@@ Appendix A Solving for the Relativistic ISO
$\left.\frac{dR(r)}{dr}\right|_{r=r_{p}}=0.$ (A2)
@@
$\Theta(\mu)=q^{2}-\mu^{2}\left[a^{2}(\xi^{2}-1)+\frac{\lambda^{2}}{1-\mu^{2}}\right]=0,$ (A3)
@@
$\cos i=\frac{L_{z}}{\sqrt{L_{z}^{2}+Q}}=\frac{\lambda}{\sqrt{\lambda^{2}+q^{2}}}$ (A4)
as the ”inclination” of the orbit. One of the turning points in Equation A3 then satisfies:
$\mu_{+}^{2}=\cos^{2}(\pi/2-i)$ (A5)
@@
$\mathscr{C},\mathscr{B}]\lambda^{2}+[\mathscr{D},\mathscr{B}]\lambda+[\mathscr{E},\mathscr{B}]=0$ (A6)
@@
$\mathscr{B}\xi^{2}+\mathscr{C}\lambda^{2}+\mathscr{D}\lambda+\mathscr{E}=0$ (A8)
@@
$q^{2}=\mu_{+}^{2}\left[a^{2}(\xi^{2}-1)\right]$ (A9)
So far, for given values of $r_{a}$, $r_{p}$, and $i$, we can solve for $\xi$, $\lambda$, and $q^{2}$ using Equations A6,  A8, and A9.
The remaining two roots of Equation A1 are given by
$r_{3}=\alpha+\sqrt{\alpha^{2}-\beta},~r_{4}=\alpha-\sqrt{\alpha^{2}-\beta}$ (A10)
@@
$\alpha=-\frac{\xi^{2}}{1-\xi^{2}}-\frac{r_{a}+r_{p}}{2},~\beta=-\frac{a^{2}q^{2}}{1-\xi^{2}}\frac{1}{r_{a}r_{p}}$ (A11)
@@
    Calculate the apocenter distance $r_{a}=r_{p}/(1-e)$. Then, solve for $\lambda$ using Equation A6.
@@
    Solve for $\xi$ and $q^{2}$ using Equations A8 and A9, respectively. Then, calculate $r_{3}$ and $r_{4}$ using Equation A10.
@@
$r_{p}=\frac{2(3+e)}{1+e}r_{g},~\lambda={\rm sgn}(L_{z})\frac{(3+e)^{3/2}}{2^{1/2}(1+e)^{1/2}}$ (A12)
Thus, in the limit $e\rightarrow 1$, we have $r_{p}\rightarrow 4r_{g}$ and $|\lambda|\rightarrow 4$. For a circular orbit ($e=0$), $r_{p}=6r_{g}$. The left panel of Figure A shows the solution for $r_{p}$ for orbits around a Schwarzschild ($a=0$) or a maximally spinning ($a=1$) MBH.
Appendix B Mapping to a Newtonian Loss Cone
@@
This mapping is not unique, however, as there is freedom in the choice of relativistic coordinates. For example, the transformation can be performed using the original BL coordinates, the local non-rotating rest frame, or harmonic coordinates. Nevertheless, all these mappings should converge at sufficiently large distances from the MBH. As the majority of particles in GNC have apocenter distances $\gg r_{g}$, we perform the mapping at the orbital apocenter under the Schwarzschild metric. In this case, the Newtonian semi-major axis ($a_{\rm NW}$, in units of $r_{g}$), eccentricity ($e_{\rm NW}$), and inclination ($i_{\rm NW}$) of the orbit are given by
$a_{\rm NW}$ $=-\frac{1}{2E_{\rm NW}},$ (B1)
$e_{\rm NW}$ $=1-v^{2}r,$
$\cos i_{\rm NW}$ $=\frac{v_{\phi}}{v}=i,$
where
$E_{\rm NW}$ $=\frac{1}{2}v^{2}-\frac{1}{r},$ (B2)
$v$ $=(v_{\theta}^{2}+v_{\phi}^{2})^{1/2},$
$v_{\theta}$ $=\pm\frac{q}{r}\left(1-\frac{2}{r}\right)^{1/2}$
$v_{\phi}$ $=\frac{\lambda}{r}\left(1-\frac{2}{r}\right)^{1/2}$
Given the values of $\lambda$ and $q$ for an ISO from the previous section, along with an inclination angle $i$, we can obtain the Newtonian pericenter $r_{\rm p,NW}=a_{\rm NW}(1-e_{\rm NW})$ using Equation B1. For nearly parabolic orbits where $e_{\rm NW}\simeq e\rightarrow 1$, the Newtonian angular momentum approaches $L_{\rm NW}=(r_{\rm p,NW}(1+e_{\rm NW}))^{1/2}\rightarrow(2r_{\rm p,NW})^{1/2}$, which in turn approaches $\lambda$. For a Schwarzschild MBH, as shown at the end of Section A, $\lambda\rightarrow 4$ when $e\rightarrow 1$. Thus, we immediately find that $r_{\rm p,NW}=8r_{g}$. This is the commonly adopted value for the pericenter of an ISO around a Schwarzschild MBH in studies using Newtonian approximations. The solutions for $r_{\rm p,NW}$ for different MBH spin parameters $a$ and orbital inclination angles are shown in Figure A.

### v2

@@ Appendix B Solving for the Relativistic ISO
$\left.\frac{dR(r)}{dr}\right|_{r=r_{p}}=0.$ (B2)
@@
$\Theta(\mu)=q^{2}-\mu^{2}\left[a^{2}(\xi^{2}-1)+\frac{\lambda^{2}}{1-\mu^{2}}\right]=0,$ (B3)
@@
$\cos i=\frac{L_{z}}{\sqrt{L_{z}^{2}+Q}}=\frac{\lambda}{\sqrt{\lambda^{2}+q^{2}}}$ (B4)
as the ”inclination” of the orbit. One of the turning points in Equation B3 then satisfies:
$\mu_{+}^{2}=\cos^{2}(\pi/2-i)$ (B5)
@@
$\mathscr{C},\mathscr{B}]\lambda^{2}+[\mathscr{D},\mathscr{B}]\lambda+[\mathscr{E},\mathscr{B}]=0$ (B6)
@@
$\mathscr{B}\xi^{2}+\mathscr{C}\lambda^{2}+\mathscr{D}\lambda+\mathscr{E}=0$ (B8)
@@
$q^{2}=\mu_{+}^{2}\left[a^{2}(\xi^{2}-1)\right]$ (B9)
So far, for given values of $r_{a}$, $r_{p}$, and $i$, we can solve for $\xi$, $\lambda$, and $q^{2}$ using Equations B6,  B8, and B9.
The remaining two roots of Equation B1 are given by
$r_{3}=\alpha+\sqrt{\alpha^{2}-\beta},~r_{4}=\alpha-\sqrt{\alpha^{2}-\beta}$ (B10)
@@
$\alpha=-\frac{\xi^{2}}{1-\xi^{2}}-\frac{r_{a}+r_{p}}{2},~\beta=-\frac{a^{2}q^{2}}{1-\xi^{2}}\frac{1}{r_{a}r_{p}}$ (B11)
@@
    Calculate the apocenter distance $r_{a}=r_{p}/(1-e)$. Then, solve for $\lambda$ using Equation B6.
@@
    Solve for $\xi$ and $q^{2}$ using Equations B8 and B9, respectively. Then, calculate $r_{3}$ and $r_{4}$ using Equation B10.
@@
$r_{p}=\frac{2(3+e)}{1+e}r_{g},~\lambda={\rm sgn}(L_{z})\frac{(3+e)^{3/2}}{2^{1/2}(1+e)^{1/2}}$ (B12)
Thus, in the limit $e\rightarrow 1$, we have $r_{p}\rightarrow 4r_{g}$ and $|\lambda|\rightarrow 4$. For a circular orbit ($e=0$), $r_{p}=6r_{g}$. The left panel of Figure B shows the solution for $r_{p}$ for orbits around a Schwarzschild ($a=0$) or a maximally spinning ($a=1$) MBH.
Appendix C Mapping to a Newtonian Loss Cone
@@
However, this mapping is not unique, as there is freedom in choosing both the relativistic coordinate system and the mapping quantities. For example, the transformation can be performed using the original BL coordinates, the local non-rotating rest frame, or harmonic coordinates. The mapping quantities can also be defined by equating the conserved orbital quantities (e.g., energy and angular momentum), or by matching specific orbital elements (e.g., periapsis or apoapsis). Nevertheless, all these mappings should converge at sufficiently large distances from the MBH. As the majority of particles in GNC have apocenter distances $\gg r_{g}$,
we perform the mapping between the Newtionan and relativistic specific orbital energy $E_{\rm NW}\rightarrow E/m-1=1/\xi-1$, $z$-component angular momentum $L_{NW,z}\rightarrow L_{z}/m=\lambda/\xi$ and total angular momentum $L_{\rm NW}\rightarrow\sqrt{L_{z}^{2}+Q}/m=\sqrt{\lambda^{2}+q^{2}}/\xi$. The mappings of orbital elements now become:
$a_{\rm NW}$ $=\frac{1}{2}\frac{\xi}{\xi-1}$ (C1)
$e_{\rm NW}$ $=\sqrt{1-j^{2}}=\sqrt{1-(\lambda^{2}+q^{2})\frac{2(\xi-1)}{\xi^{3}}}$
$\cos i_{\rm NW}$ $=\frac{L_{z}}{\sqrt{L_{z}^{2}+Q}}=\frac{\lambda}{\sqrt{\lambda^{2}+q^{2}}}=\cos i$
Given the values of $\xi$, $\lambda$ and $q$ for an ISO from the previous section, we can obtain the Newtonian pericenter $r_{\rm p,NW}=a_{\rm NW}(1-e_{\rm NW})$ using Equation C1. For nearly parabolic orbits where $e_{\rm NW}\simeq e\rightarrow 1$, the Newtonian angular momentum approaches $L_{\rm NW}=(r_{\rm p,NW}(1+e_{\rm NW}))^{1/2}\rightarrow(2r_{\rm p,NW})^{1/2}$, which in turn approaches $\sqrt{\lambda^{2}+q^{2}}$. For a Schwarzschild MBH, we can always choice coordinates such that $q=0$ and $L_{\rm NW}\rightarrow\lambda$. As shown at the end of Section B, $\lambda\rightarrow 4$ when $e\rightarrow 1$. Thus, we immediately find that $r_{\rm p,NW}=8r_{g}$. This is the commonly adopted value for the pericenter of an ISO around a Schwarzschild MBH in studies using Newtonian approximations. The solutions for $r_{\rm p,NW}$ for different MBH spin parameters $a$ and orbital inclination angles are shown in Figure B.